OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20.

OpenStudy (anonymous):

20=2(20-4)?

OpenStudy (anonymous):

Would it be 32?

OpenStudy (anonymous):

no... 20=2(x-4) 20=2x-8 2x=20+8 continue...

OpenStudy (anonymous):

14?